OpenStudy (anonymous):

if the sequence below continues according to the pattern shown, what is the sum of the first 305 terms of the sequence? -3,0,6,-3,0,6,-3,0,6...

OpenStudy (anonymous):

A) -305 B) 305 C) 610 D) -610

OpenStudy (anonymous):

sum of first 3 terms = -3 + 0 + 6 = 3 sum of first 303 terms = sum of first 3*101 = 3*101 = 303 next 2 terms are -3 and 0 so sum of 305 terms = 303 + -3 + 0 = 306

OpenStudy (anonymous):

sorry 300

OpenStudy (anonymous):

305??

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

300

OpenStudy (anonymous):

ok